\begin{thebibliography}{1}

\bibitem{Balu}
R. Balasubramanian and K. Ramachandra.
\newblock Some problems of analytic number theory III.
\newblock {\em Hardy-Ramanujan J.}, 4:13--40, 1981.

\bibitem{Bourgain}
J. Bourgain.
\newblock Decoupling, exponential sums and the Riemann zeta function. 
\newblock {\em J. Amer. Math. Soc.}, 30(1):205--224, 2017.

\bibitem{Graham}
S. W. Graham.
\newblock
Large values of the Riemann zeta-function. Topics in analytic number theory (Austin, Tex., 1982), 98--126, Univ. Texas Press, Austin, TX, 1985.

\bibitem{RHB12}
D. R. Heath-Brown.
\newblock The twelfth power moment of the Riemann zeta-function.
\newblock {\em Quart. J. Math. Oxford Ser. (2)}, 29(116):443--462, 1978.

\bibitem{RHB4}
D.R. Heath-Brown.
\newblock The fourth power moment of the Riemann zeta function.
\newblock {\em Proc. London Math. Soc. (3)}, 38(3):385--422, 1979.

\bibitem{RHBZero}
D. R. Heath-Brown.
\newblock Zero density estimates for the Riemann zeta-function and Dirichlet $L$-functions.
\newblock {\em J. London Math. Soc. (2)}, 19(2):221--232, 1979.

\bibitem{RHBMean}
D. R. Heath-Brown.
\newblock Mean values of the zeta function and divisor problems.
\newblock Recent progress in analytic number theory, Vol. 1 (Durham, 1979), pp.\ 115--119, Academic Press, London--New York, 1981.



\bibitem{Aster}
D. R. Heath-Brown.
\newblock The number of Abelian groups of order at most $x$.
\newblock {\em Ast\'{e}risque}, 198-199-200:153--163, 1991.


\bibitem{IvicB1983}
A. Ivi\'{c}.
\newblock Topics in recent zeta function theory. Publ. Math. d'Orsay, Université de Paris-Sud, Dép. de Mathématique, 1983, 272 pages.

\bibitem{IvicOmega}
A. Ivi\'{c}.
\newblock The general divisor problem.
\newblock {\em J. Number Theory}, 26:73--91, 1987.


\bibitem{Ivic91}
A. Ivi\'{c}.
\newblock{Lectures on mean values of the Riemann zeta function}. Tata Institute of Fundamental Research Lectures on Mathematics and Physics, 92. Springer-Verlag, Berlin, 1991.


\bibitem{Ivic02}
A. Ivi\'{c}.
\newblock On the moments of Hecke series at central points.
\newblock {\em Funct. Approx. Comment. Math.}, 30:49--82, 2002.

\bibitem{IvicB}
A. Ivi\'{c}.
\newblock The Riemann Zeta-function. Theory and applications. Dover Publications, Inc. Mineola, NY, 2003.




\bibitem{IvicHardy}
A. Ivi\'{c}.
\newblock The Theory of Hardy's $Z$-function.
\newblock Cambridge Tracts in Mathematics, 196. Cambridge University Press, Cambridge, 2013.

\bibitem{Ivic14}
A. Ivi\'{c}.
\newblock The mean values of the Riemann zeta-function on the critical line.
\newblock{Analytic number theory, approximation theory, and special functions,} pp.\ 3--68, Springer, New York, 2014.

%\bibitem{I99}
%A. Ivi\'{c}.
%\newblock On the error term for the fourth moment of the Riemann zeta-function.
%\newblock{\em J. London Math. Soc. (2)}, 60(1):21--32, 1999.

%\bibitem{I94}
%A. Ivi\'{c}.
%\newblock Power moments of the Riemann zeta-function over short intervals.
%\newblock{\em Arch. Math. (Basel)}, 62(5):418--424, 1994.

%\bibitem{Ivic2003}
%A. Ivi\'{c}.
%\newblock On some mean value results involving $|\zeta(1/2 + it)|$.
%\newblock{\em Bull. Cl. Sci. Math. Nat. Sci. Math.}, 28:17--29, 2003.





%\bibitem{IvicLap}
%The Laplance and Mellin transforms\ldots


\bibitem{IvicM90}
A. Ivi\'{c} and Y. Motohashi.
\newblock A note on the mean value of the zeta and $L$-functions VII.
\newblock{ \em Proc. Japan Acad. Ser. A.}, 66:150--152, 1990.


\bibitem{IM94}
A. Ivi\'{c} and Y. Motohashi.
\newblock The mean square of the error term for the fourth power moment of the zeta-function.
\newblock{ \em Proc. London Math. Soc. (3)}, 69(2):309--329, 1994.

\bibitem{IvicM95}
A. Ivi\'{c} and Y. Motohashi.
\newblock On the fourth power moment of the Riemann zeta-function.
\newblock{\em J. Number Theory}, 51(1):16--45, 1995.


\bibitem{Iwaniec}
H. Iwaniec.
\newblock Fourier coefficients of cusp forms and the Riemann zeta-function.
\newblock Seminar on Number Theory, 1979-1980 (French), Exp. No. 18, 36pp., Univ. Bordeaux I, Talence, 1980.




%\bibitem{Jutila91}
%M. Jutila.
%\newblock
%Mean value estimates for exponential sums with applications to $L$-functions.
%\newblock {\em Acta Arith.}, 57(2):93--114, 1991.


%\bibitem{JutillaM}
%M. Jutila and Y. Motohashi.
%\newblock Mean value estimates for exponential sums and $L$-functions: a spectral-theoretic approach.
%\newblock {\em J. Reine Angew. Math.}, 459:61--87, 1995.

\bibitem{KatokSarnak}
S. Katok and P. Sarnak. 
\newblock
Heegner points, cycles and Maass forms.
\newblock
{\em Israel J. Math.} 84:193--227, 1993. 

%\bibitem{LauTsang}
%Y.-K. Lau and K.~M. Tsang.
%\newblock
%Omega result for the mean square of the Riemann zeta function.
%\newblock
%{\em Manuscripta Math.} 117(3):373--381, 2005.

%\bibitem{Matsumoto}
%K. Matsumoto.
%\newblock Recent developments in the mean square theory of the Riemann zeta and other zeta-functions.
%\newblock {\em Number Theory}, pp.\ 241--286, Trends Math., Birkh\"{a}user, Basel, 2000.

\bibitem{Meurman84}
T. Meurman.
The mean twelfth power of Dirichlet $L$-functions on the critical line.
\newblock {\em Ann. Acad. Sci. Fenn. Ser. A I Math. Dissertationes}, 54:44pp., 1984.

\bibitem{MotohashiV3}
Y. Motohashi.
A note on the mean value of the zeta and $L$-functions. III.
\newblock {\em Proc. Japan Acad. Ser. A Math. Sci.}, 62(4):152--154, 1986.

\bibitem{MotohashiB}
Y. Motohashi.
\newblock Spectral Theory of the Riemann Zeta-function. Cambridge Tracts in Mathematics, 127. Cambridge University Press, Cambridge, 1997.



%\bibitem{MotohashiV2}
%Y. Motohashi.
%A note on the mean value of the zeta and $L$-functions. II.
%\newblock {\em Proc. Japan Acad. Ser. A Math. Sci.}, 61(10):313--316, 1985.

%\bibitem{MotohashiV1}
%Y. Motohashi.
%A note on the mean value of the zeta and $L$-functions. I.
%\newblock {\em Proc. Japan Acad. Ser. A Math. Sci.}, 61(7):222--224, 1985.


\bibitem{Sank}
A. Sankaranararyanan.
\newblock Higher moments of certain $L$-functions on the critical line.
\newblock{\em Lithuanian Math. J.}, 47(3), 277--310, 2007.

\bibitem{Titchmarsh}
E. C. Titchmarsh
\newblock The Theory of The Riemann Zeta-Function. Second Revised (Heath-Brown) Edition. Oxford University Press, Oxford, 1986.
\end{thebibliography}

\documentclass[12pt]{article}
\author{Neea Paloj\"{a}rvi\footnote{Supported by the Finnish Cultural Foundation.}\, and Tim Trudgian\footnote{Supported by Australian Research Council Discovery Project DP240100186.} \\
School of Science, UNSW Canberra, Australia \\
 n.palojarvi@unsw.edu.au\quad  \quad timothy.trudgian@unsw.edu.au }
  \title{On the error term of the fourth moment of the Riemann zeta-function}
\usepackage{indentfirst}
\usepackage{url}
\usepackage{enumerate}
\usepackage{amsthm}
\usepackage{amsmath,mathrsfs}
\usepackage{comment}
\usepackage{fullpage}
\usepackage{amssymb}
\usepackage{booktabs}

\newtheorem{thm}{Theorem}

\newtheorem{Lem}{Lemma}
\newtheorem{remark}{Remark}
\newtheorem{cor}{Corollary}

\usepackage[dvipsnames]{xcolor}
\usepackage{todonotes}
\begin{document}
\maketitle
\begin{abstract}
\noindent
We examine the size of $E_{2}(T)$, the error term in the asymptotic formula for $\int_{0}^{T} |\zeta(1/2 + it)|^{4}\, dt$ where $\zeta(s)$ is the Riemann zeta-function. We make improvements in the powers of $\log T$ in the known bounds for $E_{2}(T)$ and $\int_{0}^{T} E_{2}(t)^{2}\, dt$. As a consequence, we obtain small logarithmic improvements for $k$th moments where $8\leq k\leq 12$. In particular, we make a modest improvement on the 12th power moment for $\zeta(s)$.
\end{abstract}
\section{Introduction}\label{leubald}
\noindent
Consider the fourth moment of the Riemann zeta-function
\begin{equation}\label{als}
\int_{0}^{T} |\zeta(1/2 + it)|^{4} \, dt = T P_{4}(\log T) + E_{2}(T),
\end{equation}
where $P_{4}(x)$ is a quartic polynomial, the coefficients of which are known. In this paper we consider three estimates involving $E_{2}(T)$.

The first is the size of $E_{2}(T)$ itself. Heath-Brown \cite{RHB4}, proved (\ref{als}) and showed that $E_{2}(T) \ll T^{7/8 + \epsilon}$. This was improved by Ivi\'{c} and Motohashi \cite{Ivic91,IvicM90,IvicM95} to
\begin{equation}\label{fraulenist}
E_{2}(T) \ll T^{2/3} (\log T)^{C_{1}},
\end{equation}
for some constant $C_{1}$.

The second is the mean-square of $E_{2}(T)$, which was also studied by Ivi\'{c} and Motohashi, who showed in \cite{IM94} that
\begin{equation}\label{laune}
\int_{0}^{T} E_{2}(t)^{2} \, dt\ll T^{2} \log^{C_{2}} T,
\end{equation}
for some $C_{2}$.

The third is the twelfth-power moment estimate for $\zeta(s)$. Heath-Brown \cite{RHB12} proved that
\begin{equation}\label{hochzeit}
\int_{0}^{T} |\zeta(1/2 + it)|^{12}\, dt \ll T^{2} \log^{17} T.
\end{equation}
At first glance, Heath-Brown's result does not involve $E_{2}(T)$, and, indeed, his proof involves the mean-square of $\zeta(s)$ in short intervals, and not the fourth-power moment in (\ref{als}). More general methods can be used: in 2007, Sankaranarayanan \cite{Sank} proved a version of (\ref{hochzeit}) with $33$ instead of $17$.

Iwaniec \cite{Iwaniec} was the first to prove an appropriate short-interval version of (\ref{als}) and to apply this to the twelfth-power estimate --- see also \cite{IvicB,IvicHardy}. This leads to $\int_{0}^{T} |\zeta(1/2 + it)|^{12}\, dt \ll T^{2 + \epsilon}$. Ivi\'{c} and Motohashi \cite{IM94,IvicM95} were able to refine some aspects of Iwaniec's method to improve this to an estimate of the form
\begin{equation*}\label{feen}
\int_{0}^{T} |\zeta(1/2 + it)|^{12}\, dt \ll T^{2} \log^{C_{3}} T,
\end{equation*}
where
\begin{equation*}\label{oder}
C_{3} = 8 + C_{2},
\end{equation*}
with $C_{2}$ coming from (\ref{laune}).
This leads us to investigate the size of $C_{2}$ in (\ref{laune}).

Motohashi \cite{MotohashiB} proved that one can take $C_{2} = 22$, and, via similar methods, $C_{1} = 8$ in (\ref{fraulenist}). In this article we show that we may take $C_{1} = 3.5$ and $C_{2} = 9$. In fact, we show slightly more in our main results below.

\begin{thm}
\label{thm:fourthMoment}
    For any constant $C\geq 0$, we have
    \begin{equation*}
        E_{2}(T) \ll T^{2/3} \left(\log{T}\right)^{3.5}(\log\log{T})^{-C}, \quad\text{as } T\to \infty.
    \end{equation*}
\end{thm}

\begin{thm}\label{die}
Let $A \in[0,2]$. Then for any constant $C\geq 0$, we have
\begin{equation*}
 \int_{0}^{T} E_{2}(t)^{A}\, dt \ll 
 T^{1+A/2}(\log{T})^{9A/2}\cdot
 \begin{cases}
  (\log\log{T})^{2A/3}, &\text{if } 0\leq A\leq 1.5, \\
  (\log\log{T})^{-C}, &\text{if } 1.5<A\leq 2, 
 \end{cases}\quad\text{as } T\to \infty.
\end{equation*}
\end{thm}

Theorem \ref{die} immediately gives us an estimate for the $k$th moment when $8\leq k \leq 12$.
\begin{cor}
\label{cor:12moment}
Let $k \in [8,12]$ and $C \geq 0$ any constant. Then 
\begin{equation*}
\int_{0}^{T} |\zeta(1/2 + it)|^{k}\, dt \ll T^{1+\frac{k-4}{8}} (\log{T})^{\frac{11k+4}{8}}
\begin{cases}
(\log\log{T})^{\frac{k-4}{6}}, &\text{if } 8\leq k\leq 10, \\
(\log\log{T})^{-C}, &\text{if } 10<k\leq 12, 
\end{cases}
\quad\text{as } T\to \infty.
\end{equation*}
\end{cor}

\begin{remark}
    If $k=12$, then Corollary \ref{cor:12moment} gives
    \begin{equation*}
        \int_{0}^{T} |\zeta(1/2 + it)|^{12}\, dt \ll T^{2} (\log{T})^{17}(\log\log{T})^{-C}
    \end{equation*}
    for any constant $C \geq 0$. This is a slight improvement to Heath-Brown's estimate \eqref{hochzeit}.  The bound provided in \cite[Theorem 7.2]{IvicB1983} for $k \in [8,12]$ is $\ll T^{1+(k-4)/8+\varepsilon}$ for any $\varepsilon>0$. We reprove this, and are able to keep track of the logarithmic powers.
\end{remark}

\begin{remark}
    Following the same structure as in the proofs of Theorem \ref{thm:fourthMoment} and Corollary \ref{cor:12moment}, we can prove that $\int_{0}^{T} |\zeta(1/2 + it)|^{k}\, dt \ll T^{k/6+\varepsilon}$ for all $k >12$ and any $\varepsilon>0$. However, as in the proof of Theorem 7.2 in \cite{IvicB1983}, we can state something stronger, namely, that for any $\varepsilon>0$
    \begin{equation*}
    \label{eq:highMoments}
        \int_{0}^{T} |\zeta(1/2 + it)|^{k}\, dt \ll T^{1+(k-6)c+\varepsilon}, \quad \text{as } T \to \infty,
    \end{equation*}
    where $|\zeta(1/2+i t)|\ll T^{c+\varepsilon}$ when $|t|\leq T$. By Bourgain \cite[Theorem 5]{Bourgain} we can choose $c=13/84$. 
\end{remark}

Central to the so-called spectral-theoretic approach are moments of the Hecke series $H_{j}(s)$ at the central point $s=1/2$. Indeed, 
let $\{\lambda_j=\kappa_j^2+1/4, \kappa_j>0\} \cup \{0\}$ be the discrete spectrum of the non-Euclidean Laplacian acting on $SL(2, \mathbb{Z})$. Moreover,
let $\alpha_j:=|\rho_j(1)|^2(\cosh{(\pi \kappa_j)})^{-1}$, where $\rho_j(1)$ is the first Fourier coefficient of the Maass wave form corresponding to $\lambda_j$ to which the Hecke series $H_j(s)$ is attached. We note that $H_j(1/2) \geq 0$ by \cite{KatokSarnak}, and from the theory of the Selberg zeta-function (see e.g. \cite[Equation (5.49)]{Ivic91}) it follows that $|\{\kappa_j: \kappa_j \leq T\}|=T^2/12+O(T)$. For a more detailed introduction to the Hecke series, see \cite{Ivic91}.

One requires upper bounds of the form
\begin{equation}\label{gluckliche}
\sum_{\kappa_{j}\leq K} \alpha_{j} H_{j}^{3}(1/2) \ll K^{2} \log^{A_{3}} K, \quad  \sum_{\kappa_{j}\leq K} \alpha_{j} H_{j}^{4}(1/2) \ll K^{2} \log^{A_{4}} K.
\end{equation}
Motohashi's method (see also \cite{Ivic14}) gives $C_{1} \leq (11 + 4A_{3})/6$. Indeed, Motohashi proved that $A_{3} \leq 8$, whence we arrived at his value for $C_{1}\leq 43/6 < 8$. Ivi\'c \cite{Ivic02} gave asymptotic expressions for the sums in (\ref{gluckliche}), whence one can take $A_{3} = 3$ and $A_{4} = 6$. We use these improvements, and some other refinements, to prove our results in the following sections.

\section{Proof of Theorem \ref{thm:fourthMoment}}
In this section, we prove Theorem \ref{thm:fourthMoment} by estimating $E_2(T)$ with the function
\begin{equation*}
    I(T,\Delta):= \frac{1}{\Delta \sqrt{\pi}} \int_{-\infty}^\infty \left|\zeta\left(1/2+i(T+t)\right)\right|^4e^{-(t/\Delta)^2} \, dt \quad (0<\Delta<T/\log{T}).
\end{equation*}
The following lemma describes the correspondence between $I(T, \Delta)$ and the fourth moment of the Riemann zeta-function. 

\begin{Lem}
\label{lemma:I4}
Let $\varepsilon \in [1/2,1)$, $T^\varepsilon\leq \Delta \leq T\exp(-\sqrt{\log{T}})$, $T_1:=T-\Delta(\log{T})^{1/2}$, $T_2:=2T+\Delta(\log{T})^{1/2}$, $T_3:=T+2\Delta(\log{T})^{1/2}$ and $T_4:=2T-2\Delta(\log{T})^{1/2}$.
Then
\begin{equation*}
    \int_{T_3}^{T_4} I(t,\Delta) \, dt+O(\log^4{T}) \leq \int_{T}^{2T}\left|\zeta(1/2+ix)\right|^4 \, dx \leq \int_{T_1}^{T_2} I(t,\Delta) \, dt+O(\log^4{T})  \quad\text{as } T\to \infty.
\end{equation*}
\end{Lem}

\begin{proof}
We will first provide an upper bound and then a lower bound. We have
\begin{equation*}
    \int_{T_1}^{T_2} I(t,\Delta) \, dt\geq \int_{T}^{2T} \left|\zeta(1/2+ix)\right|^4\left(\frac{1}{\Delta\sqrt{\pi}}\int_{T_1}^{T_2} e^{-(t-x)^2/\Delta^2} \, dt\right) \, dx.
\end{equation*}
Let us first consider the inner integral. Changing the variable to $t-x=\Delta v$, in $T \leq t \leq 2T$, yields
\begin{align}
    \frac{1}{\Delta\sqrt{\pi}}\int_{T_1}^{T_2} e^{-(t-x)^2/\Delta^2} \, dt &= \frac{1}{\sqrt{\pi}}\int_{\frac{T-x}{\Delta}-(\log{T})^{1/2}}^{\frac{2T-x}{\Delta}+(\log{T})^{1/2}} e^{-v^2} \, dv \nonumber\\
    &=\frac{1}{\sqrt{\pi}}\int_{-\infty}^{\infty} e^{-v^2} \, dv+O\left(\int_{(\log{T})^{1/2}}^\infty e^{-v^2} \, dv\right) \nonumber\\
    &=1+O\left(e^{-\log{T}}\right) \label{eq:eBetween}.
\end{align}
Using \eqref{als}, we obtain
\begin{equation*}
    e^{-\log{T}}\int_{T}^{2T}\left|\zeta(1/2+ix)\right|^4 \, dx \ll T^{-1+1}\log^4{T}=\log^{4}{T},
\end{equation*}
which proves the upper bound.

The lower bound follows similarly using the formula
\begin{equation*}
    \int_{T_3}^{T_4} I(t,\Delta) \, dt=\left(\int_{T}^{2T}+\int_{2T}^{\infty}+\int_{-\infty}^{T} \right)\left(\left|\zeta(1/2+ix)\right|^4\left(\frac{1}{\Delta\sqrt{\pi}}\int_{T_3}^{T_4} e^{-(t-x)^2/\Delta^2} \, dt\right)\right) \, dx,
\end{equation*}
estimating this as in \eqref{eq:eBetween}, and noting that
\begin{align*}
    & \left(\int_{2T}^{\infty}+\int_{-\infty}^{T} \right)\left(\left|\zeta(1/2+ix)\right|^4\left(\frac{1}{\Delta\sqrt{\pi}}\int_{T_3}^{T_4} e^{-(t-x)^2/\Delta^2} \, dt\right)\right) \, dx \\
    &\quad\quad \ll \left(\int_{3T}^{\infty}+\int_{-\infty}^{-1}+\int_{2T}^{3T}+\int_{-1}^T \right)\left(\frac{1}{\Delta}\int_{T_3}^{T_4} |x|^{2/3} e^{-(t-x)^2/\Delta^2} \, dt\right) \, dx \\
    &\quad\quad \ll \int_{T_3}^{T_4} \Delta \left(\frac{e^{-(3T-T)^2/\Delta^2}}{T^{1/3}}+e^{-(t+1)^2/\Delta^2}\right)+T^{2/3+1}\left(e^{-(2T-t)^2/\Delta^2}+e^{-(t-T)^2/\Delta^2}\right) \, dt \\
    &\quad\quad \ll 1.
\end{align*}
\end{proof}

Now we use Lemma \ref{lemma:I4} to derive Theorem \ref{thm:fourthMoment}.

\begin{proof}[Proof of Theorem \ref{thm:fourthMoment}]
The proof follows similar ideas as in \cite[Section 5.2]{MotohashiB} and \cite{Ivic14} with some refinements. Namely, we will show that $E_2(2T)-E_2(T)$ is asymptotically smaller than the desired upper bound by using our updated Lemma \ref{lemma:I4}, cutting the sum $S(T, \Delta)$ in different points and using a new substitution for $\Delta$.

Using \eqref{als}, we have
\begin{equation*}
    E_2(2T)-E_2(T)=\int_{T}^{2T} |\zeta(1/2 + it)|^{4} \, dt+2T P_{4}(\log {(2T)}) -T P_{4}(\log T).
\end{equation*}
By Lemma \ref{lemma:I4} and \cite[(92) \& (93)]{Ivic14} for fixed $\varepsilon \in [1/2,1)$ and $T^\varepsilon \leq \Delta \leq T\exp(-\sqrt{\log{T}})$, the right-hand side is at most
\begin{equation}
\label{eq:ErrorBound}
     S\left(2T+\Delta(\log{T})^{1/2}, \Delta\right)-S\left(T-\Delta(\log{T})^{1/2}, \Delta\right)+O\left(\Delta \log^{4.5}{T}\right)+O\left(T^{1/2}\log^B{T}\right),
\end{equation}
where $B$ is a constant and 
\begin{equation*}
    S(T, \Delta):=\pi\sqrt{\frac{T}{2}}\sum_{j=1}^\infty \alpha_j\kappa_j^{-3/2}H_j^3(1/2)\cos\left(\kappa_j\log{\frac{\kappa_j}{4eT}}\right)e^{-\left(\frac{\Delta \kappa_j}{2T}\right)^2}.
\end{equation*}
Similarly, we get a lower bound where $S\left(2T+\Delta\log{T}, \Delta\right)-S\left(T-\Delta\log{T}, \Delta\right)$ is replaced by $S\left(2T-2\Delta\log{T}, \Delta\right)-S\left(T+2\Delta\log{T}, \Delta\right)$. Next, we will estimate the term $S(T, \Delta)$.

Let $A>0$ be sufficiently large. As in \cite[p. 37]{Ivic14}, the contribution from the part where $\kappa_j >AT\Delta^{-1}\sqrt{\log{T}}$ is $O(1)$. We divide the rest of the sum into parts $\kappa_j \leq AT\Delta^{-1}(\log\log{T})^{1/2}$ and $AT\Delta^{-1}(\log\log{T})^{1/2}<\kappa_j \leq AT\Delta^{-1}\sqrt{\log{T}}$. By partial summation and the first estimate in \eqref{gluckliche}, in the first case we have
\begin{align}
    &\sum_{\kappa_j \leq AT\Delta^{-1}(\log\log{T})^{1/2}} \alpha_j\kappa_j^{-3/2}H_j^3(1/2)\cos\left(\kappa_j\log{\frac{\kappa_j}{4eT}}\right)e^{-\left(\frac{\Delta \kappa_j}{2T}\right)^2} \nonumber\\
    &\quad\quad \ll \left(T\Delta^{-1}(\log\log{T})^{1/2}\right)^{1/2}\left(\log{(T\Delta^{-1})}\right)^{A_3}+\int_1^{T\Delta^{-1}(\log\log{T})^{1/2}} \frac{\log^3{u}}{u^{1/2}} \, du \nonumber\\
    &\quad\quad \ll \left(T\Delta^{-1}(\log\log{T})^{1/2}\right)^{1/2}\left(\log{(T\Delta^{-1})}\right)^{A_3} \label{eq:SFirst}.
\end{align}
Similarly, the second sum is
\begin{align}
    &\sum_{ AT\Delta^{-1}(\log\log{T})^{1/2}<\kappa_j \leq AT\Delta^{-1}\sqrt{\log{T}}} \alpha_j\kappa_j^{-3/2}H_j^3(1/2)\cos\left(\kappa_j\log{\frac{\kappa_j}{4eT}}\right)e^{-\left(\frac{\Delta \kappa_j}{2T}\right)^2} \nonumber\\
    &\quad\quad \ll \frac{\left(T\Delta^{-1}\sqrt{\log{T}}\right)^{1/2}\left(\log{(T\Delta^{-1})}\right)^3}{T^{A^2/4}}+\left(T\Delta^{-1}(\log\log{T})^{1/2}\right)^{1/2}\left(\log{(T\Delta^{-1})}\right)^{A_3-1/4} \nonumber\\
    &\quad\quad\quad+\left(\frac{\Delta}{T}\right)^2\int_{AT\Delta^{-1}(\log\log{T})^{1/2}}^{ AT\Delta^{-1}\sqrt{\log{T}}} \frac{u^{3/2}(\log{u})^{A_3}}{e^{\left(\frac{\Delta u}{2T}\right)^2}} \, du \nonumber \\
    &\quad\quad \ll \left(T\Delta^{-1}(\log\log{T})^{1/2}\right)^{1/2}\left(\log{(T\Delta^{-1})}\right)^{A_3}+\left(T\Delta^{-1}\right)^{1/2}\left(\log{T}\right)^{5/4} \label{eq:SSecond}
\end{align}
for sufficiently large $A$.
Setting $\Delta=T^{2/3}(\log{T})^B (\log\log{T})^D$ for some constants $B_1, B_2$ to \eqref{eq:SFirst} and \eqref{eq:SSecond}, we obtain
\begin{equation*}
    S(T, \Delta) \ll T^{2/3}\left((\log{T})^{A_3-B_1/2}(\log\log{T})^{1/4-B_2/2}+(\log{T})^{5/4-B_1/2}(\log\log{T})^{-B_2/2}\right).
\end{equation*}
Next, we optimize the constants $B_1$ and $B_2$.

The term $\Delta\log^{4.5}{T}$ in \eqref{eq:ErrorBound} contributes the term $T^{2/3}(\log{T})^{4.5+B_1}$. By \cite[Theorem 1]{Ivic02} we have $A_3=3$. Thus we choose $B_1=-1$ and $B_2=1/2-2C$. Combining the estimates to the bound \eqref{eq:ErrorBound}, we can conclude that $E_2(2T)-E_2(T)\ll T^{2/3} \left(\log{T}\right)^{3.5}/(\log\log{T})^C$.
\end{proof}

\begin{remark}
    Using the method above, the potential improvements for $E_2(T)$ should come improving the estimate for $S(T, \Delta).$ 
    We also note that choosing an exponent less than $1/2$ for $\log{T}$ in $T_1,\ldots, T_4$ in Lemma \ref{lemma:I4} induces an error term $O(T)$ in estimate \eqref{eq:ErrorBound}, which would be too large.
\end{remark}

\section{Proofs of Theorem \ref{die} and Corollary \ref{cor:12moment}}
We first need to estimate an integral involving the $\Gamma$-function.
\begin{Lem}
\label{lemma:gammaIntegral}
    Let $\omega:=1/\log{T}$. Then 
    \begin{equation*}
        \int_{(\omega)} \left|\Gamma(z)\right| \, dz \ll \log\log{T} \quad\text{as } T \to \infty.
    \end{equation*}
\end{Lem}

\begin{proof}
    Since $|\Gamma(z)|=|\Gamma(\overline{z})|$, it is sufficient to consider non-negative imaginary parts. By \cite[Equation (4.12.2)]{Titchmarsh}
    \begin{equation*}
        \int_1^\infty \left|\Gamma(\omega+i x)\right| \, dx=\int_1^\infty x^{\omega-1/2}e^{-\pi x/2}\left(1+O\left(\frac{1}{x}\right)\right) \, dx \ll \int_1^\infty e^{-\pi x/2} \, dx \ll 1
    \end{equation*}
    for large enough $T$. Moreover, a Laurent series expansion shows that $\Gamma(z)=1/z+O(1)$ when $|z|$ is bounded. Hence, we have
    \begin{align*}
        \int_0^1 \left|\Gamma(\omega+i x)\right| \, dx\leq \int_0^1 \frac{1}{|\omega+ix|} \, dx+O(1)\ll \log{\left(\frac{2\sqrt{\omega^2+1}+2}{\omega^2}+1\right)}+O(1) \ll \log\log{T}
    \end{align*}
    as $T \to \infty$.
\end{proof}

The idea behind the proof of Theorem \ref{die} is to show that the function $E_2(t)$ is not often very large. This is done in the following lemma.

\begin{Lem}
\label{lemma:Subintervals}
    Let $T \leq t_1 <t_2<\ldots <t_R \leq 2T$ be points such that $t_{r+1}-t_r \geq V(\log{T})^{-a}(\log\log{T})^{b}$ and $|E_2(t_r)| \geq V$ for all $r=1,\ldots, R-1$, where
    \begin{equation*}
        T^{1/2} (\log T)^{c}(\log\log{T})^b \leq V \leq T^{2/3} (\log T)^{d}
    \end{equation*}
    and $a,b,c, d$ are constants with $a>6$.
    Then
    \begin{equation}
    \label{eq:R}
        R \ll T^2V^{-3}(\log{T})^{a+9}(\log\log{T})^{1-b/2}.
    \end{equation}
\end{Lem}

\begin{proof}
Let us first choose $c(\varepsilon_1)$ is such a way that for any given small $\varepsilon_1>0$ we have $t_r 2^{-L} \asymp T^{3/4-\varepsilon_1}$, where $L:=[c(\varepsilon_1)\log{T}]$. Moreover, let us also set $\Delta:= V(\log T)^{-g}(\log\log{T})^{-b}$ for some constant $g>5$ and $g \leq a-1$. By the definitions of $R$ and $V$ and Theorem \ref{thm:fourthMoment}, we have
\begin{equation*}
    RV \leq \sum_{r=1}^R E_2(t_r) =\sum_{r=1}^R\sum_{l=1}^L \left(E_2\left(t_r 2^{1-l}\right)-E_2\left(t_r2^{-l}\right)\right)+O\left(RT^{(1-\varepsilon_1)/2}\right).
\end{equation*}
By \cite[p. 36]{Ivic14} and \cite[Equation (1.13)]{IvicM95}, the right-hand side is
\begin{align}
    &=\sum_{r=1}^R\sum_{l=1}^L \left(S(2^{1-l}t_r+\Delta\log{(t_r 2^{-l})}, \Delta)-S(2^{-l}t_r-\Delta\log{(t_r 2^{-l})}, \Delta)\right) \nonumber \\
    &\quad\quad+O\left(R\Delta \log^{5}{T}\right)+O\left(RT^{1/2}\right). \label{eq:RV}
\end{align}
Using $\Delta= V(\log T)^{-g}(\log\log{T})^{b}$, the second last term can be written as 
\begin{equation*}
    O\left(RV \log^{5-g}{T}(\log\log{T})^{b}\right).
\end{equation*}
Thus, we concentrate on estimating the sums in terms of $S$.

We set $\tau(r,l):=t_r 2^{-l}-\Delta \log{(t_r 2^{-l})}$. Similarly as in \cite[p. 317]{IM94}, we can conclude that
\begin{align*}
    & \sum_{r=1}^R S\left(\tau(r,l), \Delta\right)=2^{-1/2}\pi \sum_{r=1}^R \tau(r,l)^{1/2}\sum_{\kappa_j \leq T\Delta^{-1}\log{T}} \alpha_j \kappa_j^{-3/2} H_j(1/2)^3\cos{\left(\kappa_j\log{\frac{4 e \tau(r,l)}{\kappa_j}}\right)} \cdot \\
    &\quad\quad \cdot \frac{1}{2 \pi i} \int_{(\omega)} \Gamma(z)\left(\frac{2\tau(r,l)}{\Delta \kappa_j}\right)^{2z} \, dz +o(1),
\end{align*}
where $\omega:=1/\log{T}$. Changing the order of the summation and setting $U:=T\Delta^{-1}\log{T}$, we obtain
\begin{equation*}
    \sum_{r=1}^R S\left(\tau(r,l), \Delta\right) \ll 1+\left(1+\sum_{m=1}^{\log_2{(U/2)}}\right)\int_{(\omega)} \left|\Gamma(z)\right| \sum_{2^{-m}U \leq \kappa_j \leq 2^{1-m}U} \frac{\alpha_j H_j(1/2)^3}{(2^{-m}U)^{3/2}} \left|\sum_{r=1}^R \tau(r,l)^{1/2+2z+i\kappa_j}\right|\, dz.
\end{equation*}
We will now estimate the sum over $2^{-m}U \leq \kappa_j \leq 2^{1-m}U$ and then combine the estimates.

We use the Cauchy--Schwarz inequality to show
\begin{align}
     & \sum_{K \leq \kappa_j \leq 2K} \frac{\alpha_j H_j(1/2)^3}{K^{3/2}} \left|\sum_{r=1}^R \tau(r,l)^{1/2+2z+i\kappa_j}\right| \ll K^{-3/2}\left(\sum_{K< \kappa_j \leq 2K} \alpha_j H_j(1/2)^4\right)^{1/2} \cdot \nonumber\\
     &\quad\quad \cdot \left(\sum_{K< \kappa_j \leq 2K} \alpha_j H_j(1/2)^2 \left|\sum_{r=1}^R \tau(r,l)^{1/2+2z+i\kappa_j}\right|^2 \right)^{1/2}. \label{eq:spectral2}
\end{align}
By the second estimate in (\ref{gluckliche}) we get
\begin{equation}
\label{eq:Hj4Sum}
    \sum_{K< \kappa_j \leq 2K} \alpha_j H_j(1/2)^4 \ll K^2\log^6{K}.
\end{equation}
Moreover, since by assumptions $|t_{r+1}-t_r| \geq V(\log{T})^{-a}(\log\log{T})^b$ and $\left|\tau(r,l)^{1/2+2z+i\kappa_j}\right|\ll 2^{-l/2}T^{1/2}$, we can apply \cite[Theorem 3.1]{MotohashiB}, and get
\begin{equation}
\label{eq:Hj2Sum}
    \sum_{K< \kappa_j \leq 2K} \alpha_j H_j(1/2)^2 \left|\sum_{r=1}^R \tau(r,l)^{1/2+2z+i\kappa_j}\right|^2  \ll R(K+TV^{-1}(\log{T})^{a}(\log\log{T})^{-b})2^{-l}TK\log{K}.
\end{equation}
Hence, for $K \ll T\Delta^{-1}\log{T} \ll TV^{-1}(\log{T})^a(\log\log{T})^{-b}$, the right-hand side of \eqref{eq:spectral2} is
\begin{equation*}
    \ll 1+2^{-l/2}R^{1/2}TV^{-1/2}(\log{T})^{a/2}(\log\log{T})^{-b/2}\sum_{m=1}^{\log_2{(U/2)}} (\log{(2^{-m}U)})^{3.5}\int_{(\omega)} \left|\Gamma(z)\right| \, dz.
\end{equation*}
Since $\log{(2^{-m}U)}\asymp \log{T}$ for all $m$, by Lemma \ref{lemma:gammaIntegral} the right-hand side is
\begin{equation*}
    \ll  1+2^{-l/2}R^{1/2}TV^{-1/2}(\log{T})^{a/2+4.5}(\log\log{T})^{1-b/2}.
\end{equation*}
Hence, keeping in mind that $a-1\geq g > 5$, the claim follows using estimate \eqref{eq:RV}.
\end{proof}

\begin{proof}[Proof of Theorem \ref{die}]
If
\begin{equation*}
    |E_2(t)| \leq T^{1/2}(\log{T})^{9/2}(\log\log{T})^{b},
\end{equation*}
where $b=2/3$ if $A\leq 1.5$ and $b=\frac{-1-C}{A-1.5}$ if $A > 1.5$, then the claim clearly holds. Hence, let us assume that this is not the case. We shall divide the integral from $0$ to $T$ into sub-integrals from $2^{-m}T$ to $2^{1-m}T$ for $m=1,2,\ldots$. We will consider a more general case and hence for each $m$, let $\mathfrak{I}_m$ denote the subset of $[2^{-m}T,2^{1-m}T]$ where 
$$ 
|E_2(t)| \geq (2^{-m} T)^{1/2}\cdot
(\log{(2^{-m}T)})^{9/2}(\log\log{(2^{-m}T)})^{b}. 
$$
We note that the upper bound for $|E_2(t)|$ is provided by Theorem \ref{thm:fourthMoment}. Moreover, we divide each interval $[2^{-m}T,2^{1-m}T]$ into sub-intervals of length $V(\log{(2^{-m}T)})^{-7}(\log\log{(2^{-m}T)})^{b}$. By Lemma \ref{lemma:Subintervals} we obtain
\begin{align*}
& \int_{0}^{T} E_{2}(t)^{2}\, dt \ll \sum_{m=1}^{\infty} \int_{\mathfrak{I}_m(V)} E_{2}(t)^{2}\, dt \\
 &\quad\quad\ll \sum_{m=1}^{\infty} V^A\cdot V(\log{(2^{-m}T)})^{-7}(\log\log{(2^{-m}T)})^{b}(2^{-m}T)^2V^{-3}(\log{(2^{-m}T)})^{7+9}(\log\log{T})^{1-b/2} \\
 &\quad\quad \ll T^{1+A/2} (\log{T})^{9A/2}(\log\log{T})^{1+(A-1.5)b}.
\end{align*}
We obtain the result by substituting the value of $b$.
\end{proof}

We note that it is easy to use Theorem \ref{die} to produce Corollary \ref{cor:12moment}.

\begin{proof}[Proof of Corollary \ref{cor:12moment}]
    The claim follows from \cite[Theorem 4]{IM94} with $k=2$ and $A=(k-4)/4$, and hence by Theorem \ref{thm:fourthMoment} with $c(k,A)=A/2$ and $d(k,A)=9A/2$. Note that we have to multiply the last term in \cite[(2.4)]{IM94} with $\log\log{T}$ (with the suitable power) because of the respective factor in Theorem \ref{die}. This gives us the extra $\log\log{T}$ term in Corollary \ref{cor:12moment}.
\end{proof}

\section{Discussion, extension, and future work}
It may be possible to make further improvements to Heath-Brown's 12th moment estimate in \eqref{hochzeit} by improving the $(\log{T})^9$ in \eqref{eq:R} or improving the result in Theorem 4 in \cite{IM94}. We identify the following avenues for potential improvement.
\begin{itemize}
    \item Is it possible to improve estimate \eqref{eq:Hj2Sum}? Note that the bound in \eqref{eq:Hj4Sum} is optimal.
    \item In \cite[Equation (4.5)]{IM94}, there is the estimate
    \begin{equation*}
        \left|\zeta(1/2+it)\right|^{2k} \ll (\log{t})\left(1+\int_{t-1/3}^{t+1/3} \left|\zeta(1/2+it + iu)\right|^{2k} \, du \right).
    \end{equation*}
    It is not clear to us that the $\log{t}$ term, which comes from growth estimates of $\zeta(1/2 + it)$, can be reduced. 
\end{itemize}

We note that Heath-Brown \cite{RHBMean} used the 12th power moment to prove $$\int_{0}^{T} |\zeta(5/8 + it)|^{8}\, dt \ll T (\log T)^{38},$$ and in \cite{RHBZero} to improve on certain zero-density results. We expect that these results could be improved very slightly using our work. Similarly, Graham \cite{Graham} (see also \cite[p.\ 179]{Titchmarsh}) adapted Heath-Brown's method to prove $$\int_{0}^{T} |\zeta(5/7 + it)|^{196}\, dt \ll T^{14} (\log T)^{425}.$$ We note (very briefly) the 12th power moment for Dirichlet $L$-functions, also with logarithmic powers, in Motohashi \cite{MotohashiV3}, improving work by Meurman \cite{Meurman84}.

While we have not pursued potential improvements to these results, we do list one final result indirectly related to this work. 

Let $a(n)$ denote the number of non-isomorphic Abelian groups of order $n$, and let $A(x) = \sum_{n\leq x} a(n)$. The problem is to estimate $\Delta(x)$, the error term in the approximation $A(x) = \sum_{k=1}^{5}c_{k}x^{1/k} + \Delta(x)$.

Ivi\'{c} \cite{IvicOmega} proved $\int_{1}^{X} \Delta(x)^{2} \, dx = \Omega(X^{4/3} \log X)$ --- see also \cite{Balu}. Heath-Brown \cite{Aster} gave the following upper bound that is sharp up to logarithmic powers
\begin{equation}\label{barge}
\int_{1}^{X} \Delta(x)^{2} \, dx \ll X^{4/3} (\log X)^{89}.
\end{equation}
One of the ingredients in this proof is indeed the 12th power moment. Initially we had hoped for a larger improvement in our Corollary \ref{cor:12moment}, and hence a chance to improve on this result. Indeed, since Heath-Brown noted that he had `made no attempt to obtain a good exponent for the power of $\log X$' in (\ref{barge}), we had hoped that a careful analysis of the role of the 12th power moment could reduce the power of 89 in (\ref{barge}) substantially. 

Heath-Brown's proof requires careful estimation of mean values of various Dirichlet polynomials. This analysis is done by splitting the ranges of summation into four cases and by frequently using the following fact. If $A_{i}, \alpha_{i} \geq 0$ and $\sum \alpha_{i} = 1$, then $\min\{A_{1}\, \ldots, A_{k}\} \leq A_{1}^{\alpha_{1}}\cdots A_{k}^{\alpha_{k}}$. Where powers of $\log X$ are concerned, Heath-Brown simply bounds the minimum of the $A_{i}$'s by the \textit{maximum} of the $A_{i}$'s. By retaining the exponents at each step we make a slight, cheap improvement. So, for example, the exponent of 65 in \cite[(3.1)]{Aster} can be reduced to $77/2$. This actually comes from Heath-Brown's `Case 1' (the only case not to invoke the 12th power moment), and is the largest of the four estimates. A quick examination of the proof shows that
\begin{equation}\label{barge2}
\int_{1}^{X} \Delta(x)^{2} \, dx \ll X^{4/3} (\log X)^{62.5}.
\end{equation}
Splitting into slightly different cases could bring 12th power moment estimate to the fore, in which case future improvements of Corollary \ref{cor:12moment}, could reduce the estimate in (\ref{barge2}) still further.


\section*{Acknowledgements}
We are grateful to Aleksander Simoni\v{c} for the discussions concerning the manuscript. 


 \end{document}